OpenStudy (anonymous):

Identify all zeros of the polynomial function f(x)=x63-3x^2+4x-12

OpenStudy (anonymous):

* x^3

OpenStudy (anonymous):

not x63

OpenStudy (campbell_st):

ok... looks like x = 3 is a zero.. f(3) = 0 when you substitute so group in pairs \[f(x) = x^2 + 4x - 3x^2 - 12\] which becomes \[f(x) = x(x^2 +4) -3(x^2 + 4)\] which factors to \[f(x) = (x -3)(x^2 + 4)\]

OpenStudy (campbell_st):

hope this makes sense...

OpenStudy (campbell_st):

oops should read \[f(x) = x^3 + 4x - 3x^2 - 12\]

OpenStudy (anonymous):

Im not sure what the answer should look like I got -1,2,-2

OpenStudy (campbell_st):

well here is my point, if a polynomial has a zero at x = -1 then f(-1) = 0 so looking at your solutions f(-1) = (-1)^3 - 3(-1)^2 + 4(-1) - 12 f(-1) = -1 -3 - 4 -12 f(-1) = -20 x = -1 can't be a zero f(-2) = (-2)^3 - 3(-2)^2 + 4(-2) - 12 = -8 -12 - 8 - 12 = -60 so x = -2 can't be a zero f(2) = (2)^3 - 3(2)^2 + 4(2) - 12 = 8 - 12 + 8 - 12 f(2) = -8 so x = 2 isn't a zero so perhaps you are looking at the wrong answer.

OpenStudy (anonymous):

ok-3 and -4?

OpenStudy (campbell_st):

here is a graph of your function

OpenStudy (campbell_st):

and - 4 won't work either \[f(x) = (x -3)(x^2 + 4)\] the zeros occur when x - 3 = 0..... and/or \[x^2 + 4 = 0\] so x = 3 is easy to find as a zero by now you have the problem of \[x^2 = -4\] you are now dealing with complex numbers as the square root of -4 is undefined in the real number system

OpenStudy (campbell_st):

so you function only has 1 real root, x = 3

OpenStudy (campbell_st):

hope this all makes sense.

OpenStudy (anonymous):

I get it, i just did it all over again thank you so much!